satellite73 (satellite73):

\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]\[-\frac{b}{2a}\]

OpenStudy (anonymous):

spam >.>

OpenStudy (anonymous):

i repeat \[-\frac{b}{2a}\]

OpenStudy (anonymous):

SPAM >.> lolol

OpenStudy (anonymous):

\[\huge{-\frac{b}{2a}}\]

OpenStudy (anonymous):

2ax +b

OpenStudy (anonymous):

spams <.<

OpenStudy (anonymous):

\[\color{red}{\text{minus bee over two ay}}\]

OpenStudy (anonymous):

5 symbols minus b over 2 a

OpenStudy (anonymous):

i refuse to acknowledge!

OpenStudy (anonymous):

7 syllables min us b o ver 2 a

OpenStudy (anonymous):

I guess this is a rant..?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

he started it >.>

OpenStudy (anonymous):

some poor student asks "how to i find the vertex of ..."

OpenStudy (anonymous):

and the next thing you know all hell breaks loose

OpenStudy (anonymous):

half way between the zeros!

OpenStudy (anonymous):

<.< i dont think it was that dramtic >.> lolol

OpenStudy (anonymous):

take the derivative!

OpenStudy (anonymous):

complete the square!

OpenStudy (anonymous):

blah blah blah all this work. keep track of this, keep track of that

OpenStudy (anonymous):

I'm writing all these down, go slower...

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

i like the "halfway between the zeros" one. im gonna use that.

OpenStudy (anonymous):

no one has a problem with \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] but if you erase the radical they have a fit!

OpenStudy (anonymous):

I kinda like "halfway between the zeros" as well..

OpenStudy (anonymous):

i am going to get a tattoo.

OpenStudy (anonymous):

satelitte , get some beer and relax

OpenStudy (anonymous):

i was thinking about getting a math tatoo sometime. I dont know what i would get though O.o

OpenStudy (anonymous):

these poor people as soon as i see that question "find the vertex..." i know they are in for a bumpy ride

OpenStudy (anonymous):

Pi seems a little too cliche

OpenStudy (anonymous):

put value of pi to 100 digits

OpenStudy (anonymous):

even the otherwise sane and probably quite lovely myininaya starts in on some incompressible nonsense

OpenStudy (anonymous):

that tatoo is awesome o.O i dont think i would put it somewhere so visible though <.<

OpenStudy (anonymous):

for the last time i say it: FIRST COORDINATE IS MINUS B OVER 2 A. SECOND COORDINATE IS WHATEVER YOU GET.

OpenStudy (anonymous):

it is nice. i am going to get the same one, but without the radical part

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

so we can all have a fit? haha

OpenStudy (anonymous):

btw i did not make up that "half way between the zeros" part. someone actually tried to show some poor lost soul how to find the vertex that way

OpenStudy (anonymous):

Hope that doesn't get me moderated....

OpenStudy (anonymous):

whoa